OpenStudy (anonymous):

using your function explain to martians how to slove for F(3). show your work and explain each step using complete senctens

OpenStudy (anonymous):

What Do You Think

OpenStudy (anonymous):

Using your function, explain to the Martians how to solve for f(3). Show your work and explain each step using complete sentences.f(x)= -0.5x+1 f(x)=3x-8f(3)= -0.5(3)+1 f(3)=3(3)-8f(3)= -1.5+1 f(3)=9-8f(3)= -0.5 f(3)=1For the first function, you first multiply -0.5 and 3 then you add -1.5 and 1 to get your answer. For the second function first you multiply 3 and 3, then you subtract 9 and 8 to get your answer.

OpenStudy (anonymous):

http://openstudy.com/study#/updates/530913e2e4b0e03b610f14ed Check This Out

OpenStudy (anonymous):

Medal?